# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v1

@@ 8 Acknowledgments
The authors acknowledge the benevolent support of the Department of Science and Technology (DST), Government of India. We thank Dr. Heloise Stevance for constructive suggestions and helpful comments on this work. This research has made use of the High Performance Computing (HPC) resources  [^2] https://www.iiap.res.in/centers/main-campus/computing-and-i-t/ maintained by Anish Parwage, Engineer D and Head of the Computer & IT Infrastructure Division at the Indian Institute of Astrophysics (IIA), Bangalore, whose continuous support and assistance with HPC-related issues are gratefully acknowledged.

### v2

@@ 8 Acknowledgments
The authors acknowledge the benevolent support of the Department of Science and Technology (DST), Government of India. We thank Dr. Heloise Stevance for constructive suggestions and helpful comments on this work and Mr. Raphael Baer-Way for private communication and sharing data from R. Baer-Way et al. (2025). This research has made use of the High Performance Computing (HPC) resources  [^2] https://www.iiap.res.in/centers/main-campus/computing-and-i-t/ made available by the Computer Center of the Indian Institute of Astrophysics (IIA), Bangalore.
Appendix A Choice of $\beta$ parameter for accelerated wind
As described in Section 4.3.1, due to lack of a better understanding of the driving mechanism for the RSG wind out to larger distances $\sim$ few stellar radii where dust can form to drag the gas along, we choose to parameterize this kinematic process with the $\beta$-velocity law as given by Equation (4). The normalized velocity profiles using this velocity law for various choices of $\beta$ commonly used in the literature is shown in Figure 11 which illustrates the change in the shape of the velocity profile, from convex to concave, very close to the stellar surface i.e. at $r\simeq R_{*}$, with increasing values of $\beta=0.7-4$. Comparing the shape of the velocity curves with that of observations of $\lambda$-Veloris (P. D. Bennett, 2010), we select $\beta=1.2$ for the purpose of this study, as lower values of $\beta$ result in steeper velocity profiles inconsistent with the observed values, while higher values of $\beta$ do not match the shape of the observed velocity profiles close to the stellar surface.
Figure 11: The radial variation of the wind velocity, $v(r)$, normalized to the terminal wind speed, $v_{\infty}=30~\mathrm{km~s}^{-1}$ of the RSG $\lambda$-Veloris, as per the $\beta$-law given by Equation (4) for $\beta=0.7-4$, with $v_{o}=0.1~\mathrm{km~s}^{-1}~\mathrm{and~using}~R_{*}=1.46\times 10^{13}~\mathrm{cm}$ for the radius of $\lambda$-Veloris (G. Rau et al., 2018), with its observed velocity profile given by the dotted gray line (P. D. Bennett, 2010).
Appendix B Uncertainty estimates for calculating CSM densities
The uncertainties in the values of CSM densities inferred from X-ray mass loss rates and terminal wind velocities of SNe 2023ixf (W. V. Jacobson-Galán et al., 2025) and 2020ywx (R. Baer-Way et al., 2025), according to Equation (5), have been calculated by propagating the individual uncertainties in the estimation of these two properties from measurements of X-ray fluxes and fitting of spectral features, respectively. There are also additional uncertainties that involve assumptions regarding degree of ionization in the shock propagating through the CSM post explosion as well as the composition of the CSM but a quantitative estimate of these uncertainties are beyond the scope of this work.